OpenStudy (anonymous):

solve 2 tan ^2 x + 7 tan x -15 = 0

OpenStudy (anonymous):

For \[2\tan ^{2}x + 7tanx - 15 = 0\] Solve \[2\tan ^{2}\] and \[7tanx\] First. So we get: 0.00122x + 0.12278x + -15 = 0 ==================================== Simplifying 0.00122x + 0.12278x + -15 = 0 Reorder the terms: -15 + 0.00122x + 0.12278x = 0 Combine like terms: 0.00122x + 0.12278x = 0.124x -15 + 0.124x = 0 Solving -15 + 0.124x = 0 To solve for variable 'x', move all terms containing x to the left, all other terms to the right. 0.124x = 15 Divide each side by '0.124'. The exact value of 'x' is: x = 120.9677419 Simplifying x = 121